OpenStudy (anonymous):

just need this answer real quick...timed test..got 13 minutes left: A +48.26 nC point charge is placed on the x axis at 3.007 m, and a -21.49 nC point charge is placed on the y axis at y= -6.779 m. What is the direction of the electric field at the origin?

OpenStudy (anonymous):

counterclockwise from the positive x axis

OpenStudy (anonymous):

plz tell me u can do this?

OpenStudy (ash2326):

yeah let me try

OpenStudy (ash2326):



OpenStudy (anonymous):

i got 8 minutes to submit answer

OpenStudy (anonymous):

i hate timed tests

OpenStudy (anonymous):

Just think about it yourself. First hint: You can regard the resulting electrical field as the sum of the two electrical fields of each charge.

OpenStudy (anonymous):

chris i don't have the time...i got 7 minutes

OpenStudy (anonymous):

Enough time to think ...

OpenStudy (anonymous):

afterwards i'll worry about how's and why's...right now i just need the correct answer

OpenStudy (anonymous):

26.77 is net electric charge

OpenStudy (ash2326):

Electric field at origin due to positive charge=9*10^9*48.26*10^-9/(3.007)^2( -i) = 48.0355 (-i) (toward left ) Electric field at origin due to negative charge=9*10^9*(-21.49)*10^-9/(6.779)^2 (j) =4.208(-j) towards downward direction I think clockwise from x direction

OpenStudy (anonymous):

Second hint: Use coulomb's law to calculate the contribution of each charge at the origin.

OpenStudy (anonymous):

plz for the love of god just gimme the answer and i'll work through it after i submit it

OpenStudy (anonymous):

plz

OpenStudy (ash2326):

got it?

OpenStudy (anonymous):

I'm not going to support you to cheat, sorry.

OpenStudy (anonymous):

then leave

OpenStudy (anonymous):

ashe no i don't

OpenStudy (ash2326):

are there options

OpenStudy (anonymous):

Damn, just add these numbers.

OpenStudy (anonymous):

what numbers?

OpenStudy (anonymous):

i got 2 minutes left?

OpenStudy (anonymous):

u wnana throw an airball up?

OpenStudy (ash2326):

net electric field= \[\sqrt{{48.0355}^2+{4.028}^2}\]=48.204 Volt/meter in clockwise direction from =ve x-axis

OpenStudy (ash2326):

it'd be +ve x -axis

OpenStudy (anonymous):

i submitted 128...i think it's 132 though

OpenStudy (anonymous):

thx for your help sir

OpenStudy (ash2326):

but from where you got 128

OpenStudy (anonymous):

i appreciate the effort

OpenStudy (anonymous):

i added the two numbers u had earlier....as per what chris said

OpenStudy (anonymous):

then just subtracted taht from 180

OpenStudy (anonymous):

to get the angle counterclockwise

OpenStudy (ash2326):

angle is -175 degrees

OpenStudy (ash2326):

that was magnitude

OpenStudy (anonymous):

If ash's numbers are correct it's just \[\vec E = ({-48.0},{-4.2} )^T \frac{\textrm{N}}{\textrm{C}}\], i.e. 180.5° (counter-clockwise).

OpenStudy (anonymous):

yea thx....coulda used that 15 minutes ago

OpenStudy (anonymous):

good lookin out

OpenStudy (xishem):

In the future you might consider: a.) Not wait until the last minute to do a test like this. b.) Being respectful to those who helped you. They aren't required to do so.

OpenStudy (anonymous):

to xishem....i did not wait until the last minute to do this test. It's a timed 20 minute 1 question test. I'm very respectful of those who help me. I help many many people on here in the math section. I was very respectful and appreciative of ash for taking the time to help me. Chris was not helping, therefore at the time, was not appreciative